### v1

# Superconductivity in spin-orbit coupled BaBi3 formed by in situ reduction of bismuthate films

Abstract
Oxygen-scavenging at oxide heterointerfaces has emerged as a powerful route for stabilizing metastable phases that exhibit interesting phenomena, including high-mobility two-dimensional electron gases and high Tc superconductivity. We investigate structural and chemical interactions at the heterointerface formed between Eu metal and the charged-ordered insulator, BaBiO3, which leads to emergent superconductivity at 6 K. A combination of X-ray diffraction and electron microscopy measurements shows that oxygen scavenging by the Eu adlayer leads to the formation of EuOx and the superconducting intermetallic BaBi3. The superconducting state is quasi-two-dimensional, as evidenced by the angle-dependent magnetotransport measurements and current-voltage characteristics. An in-plane upper critical field Bc2 significantly exceeding the Pauli paramagnetic limit is observed, possibly reflecting features of Ising superconductivity. The strong spin-orbit coupling at the Bi sites may pave the way for the realization of high-Tc topological superconductivity.
I Introduction
The pursuit of new materials exhibiting high-temperature superconductivity remains a vibrant and active area of research. Following the discovery of high Tc superconductivity in cuprates, attention expanded towards the other oxide systems, including bismuthates[1, 2, 3, 4]. Notably, superconductivity was also identified in potassium (K) and lead (Pb) doped BaBiO3 with a superconducting transition temperature (Tc) reaching up to 32 K for BaxKyBiO3 [3, 2, 5]. In contrast to high-Tc cuprates, where conduction is confined to two-dimensional CuO2 planes, the BiO2 planes in BaBiO3 form a three-dimensional network of corner-sharing octahedra, as shown in Figure 1A. The parent compound, monoclinic perovskite-BaBiO3 (p-BBO), is known to be a charge density wave insulator with an associated structural breathing mode distortion of the oxygen octahedral cage [6, 7, 8]. Upon doping with K or Pb, this breathing mode is suppressed, leading to the emergence of metallicity and superconductivity [2, 3, 5, 7].
Interest in Bi-rich superconductors continues due to the potential role of the strong spin-orbit coupling of Bi-6p electrons in enhancing superconductivity, and magnetic field-induced transitions from conventional s-wave to unconventional nodal d-wave superconductivity [9, 10]. Intermetallic ABi3 compounds (where A = Ba or Sr) have received attention because they exhibit superconductivity enhanced by spin-orbit coupling[11, 12, 9, 13]. Theoretical studies show that spin-orbit coupling suppresses strong Fermi nesting and increases electron-phonon coupling matrix elements associated with superconductivity[9]. At ambient pressure, BaBi3 crystallizes in a tetragonal structure (P4/mmm, a=5.06 , c=5.13 )[13, 12, 14, 15], featuring corner-sharing Bi6 octahedra (Figure 1B). Magnetotransport measurements on single crystals of BaBi3 and SrBi3 reveal superconducting transitions with onset temperatures of approximately 5.9 K and 5.6 K, respectively [12, 13]. Furthermore, scanning tunneling microscopy (STM) studies demonstrate the effect of symmetry breaking and disorder on the superconducting state in BaBi3 epitaxial films, grown in situ by molecular beam epitaxy[16]. However, to date, there have been no reports on the magnetotransport properties of thin BaBi3, due to its high air sensitivity [13].
Here, we report the emergence of superconductivity at approximately 6 K in nominally Eu-capped p-BBO thin films grown via molecular beam epitaxy. X-ray diffraction measurements reveal that oxygen scavenging by the in-vacuo deposited metallic Eu overlayer leads to the reduction of the monoclinic p-BBO layers into tetragonal superconducting BaBi3 and rock-salt BaO. Several measurements, including angle-dependent magnetotransport and current-voltage characteristics, confirm the quasi-two-dimensional nature of the superconductivity. An in-plane upper critical field Bc2 far exceeding the Pauli paramagnetic limit, raising intriguing possibilities of spin-protected superconducting states, such as Ising superconductivity[17]. These findings highlight the potential of utilizing in situ reduction of bismuthate heterostructures as a platform for exploring novel quantum states. The strong spin-orbit coupling at Bi sites, may pave the way for the realization of high-temperature topological superconductivity, a long-sought phenomenon in perovskite-based systems[18, 19].
Results and Discussion
Synthesis:
The samples were fabricated in a custom-oxide molecular beam epitaxy system. Bi, Ba, Eu, and Al were evaporated from effusion cells, with their growth rates calibrated by a quartz crystal monitor (QCM). p-BBO thin films of thicknesses ranging from 12 nm (30 unit cells) to 36 nm (90 unit cells) were grown on (001)-oriented SrTiO3 (STO) substrates at a substrate temperature of 585∘ C in an oxygen plasma with an oxygen partial pressure of 6.5x10-6 Torr and a plasma power of 250 W. Sharp reflection high-energy electron diffraction (RHEED) diffraction streaks observed during p-BBO growth shown in Figure 2A, are indicative of atomically flat p-BBO layers. The in-plane lattice parameter determined from the RHEED spacings is 4.35 , in agreement with bulk stoichiometric p-BBO [5]. After growth, the p-BBO films were cooled to 500∘ C for the deposition of Eu metal in vacuo. RHEED patterns measured during Eu growth (Figure 2A) are consistent with the formation of cubic EuO from the topotactic reduction of the p-BBO layer [20, 21, 22]. The films were capped with 3 nm AlOx at room temperature to prevent oxidation when exposed to ambient conditions.
X-ray Diffraction (XRD) Measurements
Figure 2B shows a comparison of specular scans along the (00L) direction for the nominal Eu/p-BBO sample and a single-component p-BBO film on STO. For the single-component p-BBO film, Bragg peaks are observed for the STO substrate and stoichiometric p-BBO. The calculated out-of-plane lattice constant for the p-BBO film is determined to be 4.36 in agreement with the measured values for the monoclinic bulk phase [5]. In contrast to the single-component p-BBO film, no p-BBO Bragg peaks were observed for the nominal Eu/p-BBO sample. Bragg peaks are observed at 32.316o and 34.997o corresponding to the (002) Bragg peaks of rock-salt BaO and face-centered tetragonal BaBi3, respectively. The calculated out-of-plane lattice constants are 5.52 and 5.11 in good agreement with the bulk lattice constants for BaO (5.53 ) and BaBi3 (5.13 ) [23, 12]. The lattice constant of NaCl-type EuO is 5.1439 ,[24] hence, the EuO (002) Bragg peak will overlap with the BaBi3 (002) peak. To confirm the formation of BaBi3, we deposit Al metal directly on a p-BBO film at 500∘C. We find that the Al metal effectively reduces the p-BBO film to form BaBi3 and BaO as evidenced by XRD measurements in Figure 2B.
Scanning transmission electron microscopy analysis
Scanning transmission electron microscopy (STEM) measurements were performed on the films to further investigate the structural properties of the heterostructures. Figure 2C shows a high-angle annular dark field (HAADF) STEM image of a cross-section of a nominal Eu/p-BBO film on STO. Sharp interfaces are observed between the nominal layers, indicative of insignificant intermixing. The energy-dispersive X-ray spectroscopy (EDX) profiles of the section are shown in Figure 2D. The EDX maps reveal significant chemical segregation in the nominal p-BBO region. The nominal Eu and STO regions are oxygen-rich, while the nominal p-BBO section is oxygen-poor, consistent with the reduction of p-BBO by the Eu adlayer. Ba-rich regions are observed at the nominal p-BBO/STO and Eu/p-BBO interfaces. In the nominal p-BBO region, we observe lateral phase-separated Ba and Bi-rich regions, as shown in Figure S1. The oxygen content in the Ba-rich regions is significantly higher than in the Bi-rich sections, consistent with the BaO and BaBi3 Bragg peaks observed in the XRD measurements.
Transport Measurements:
Next, we investigate the transport properties of nominal Eu/p-BBO heterostructures from 1.8 - 300 K. Figure 3A shows the temperature-dependent resistance of a nominal Eu (16 nm)/p-BBO (36 nm) heterostructure (Sample 1). As the temperature decreases to 7 K, the resistance decreases, exhibiting characteristic metallic behavior. Upon further cooling, the sample undergoes a transition to a superconducting state with a narrow transition width of approximately 0.133 K, reaching a zero resistance state at 5.77 K (). The onset superconducting temperature (defined as the temperature at which the resistance drops to 90% of its normal state value), is 6.03 K, significantly higher than the previously reported value of 3.5 K for EuO/BaBiO3 heterostructures[20] but in agreement with the reported values for BaBi3[12, 13, 9, 14, 15]. Above Tc, the resistance exhibits a linear temperature dependence, a feature not discussed in bulk BaBi3 samples. This linearity is more clearly demonstrated by the linear fitting of the resistance vs. temperature data shown in Figure S2. Such behavior is indicative of strange-metal characteristics commonly observed in unconventional superconductors [25]. The inset of Figure 3A shows the temperature-dependent resistance of sample 1 with a magnetic field (), applied out of the plane. is suppressed as increases, further confirming superconductivity. Magnetoresistance curves with at different temperatures close to the superconducting transition are plotted in Figure 3B. In the normal state above Tc, the resistance exhibits a quadratic dependence on the applied magnetic field, similar to conventional metals[26]. In particular, below Tc, a resistance peak appears above the critical magnetic field, suggesting disorder arising from granularity[27].
The temperature-dependent Hall resistance exhibits a linear dependence on the magnetic field, B, as shown in Inset Figure 3C. This linear behavior confirms the presence of a single type of charge carrier. However, a change in slope and, consequently, the sign of the Hall coefficient occurs at 65 K, which can be seen more clearly in Figure 3C. This indicates a transition from low-temperature electronic conduction to hole-dominant conduction at high temperatures. The sheet carrier density and mobility at 10 K are 1.6 1016 cm-2 and 1.2 cm2/Vs, respectively. These carrier density values are consistent with previous observations[20]. The carrier concentrations are 3-4 orders of magnitude higher than observed for the two-dimensional electron gas at SrTiO3 [28, 29] and KTaO3 [30, 31, 32, 33] interfaces, suggesting that the carriers may extend through the film thickness (). Furthermore, Figure 3D compares the resistance as a function of the nominal p-BBO thickness. The number of Eu monolayers is tuned to match the p-BBO thicknesses. The transition temperatures are comparable and independent of the nominal p-BBO thicknesses.
To further investigate the dimensionality of the system, magnetotransport measurements were performed on a second Eu/p-BBO sample (Sample 2). The zero-field temperature-dependent resistance is presented in Figure 4A, with the onset of superconductivity, , observed at 5.76 K. Temperature-dependent magnetoresistance measurements were carried out with the magnetic field applied in the out-of-plane direction perpendicular to the current direction, (Figure 4B), or in-plane direction parallel to the current direction, (Figure 4C). significantly suppresses superconductivity, with the transition nearly vanishing at B = 4 T. In contrast, for , superconductivity persists at B = 8 T, as shown in Figure 4C. The anisotropic dependence of on the direction of the magnetic field suggests quasi-two-dimensional superconductivity.
The critical fields for both out-of-plane, , and in-plane configurations were determined using the 90% normal resistance criterion. and are plotted as a function of temperature in Figure 4D. To determine the upper critical fields at T = 0 K, the data were fitted using the two-dimensional Ginzburg-Landau (GL) model for the out-of-plane configuration [34]:
where represent the extrapolated critical fields at 0 K. The extracted value of is 6.24 T, exceeding the reported value of 2.2 T for BaBi3 single crystals [12, 35]. The temperature variation of the in-plane upper critical field deviates from the 2D Ginzburg-Landau (GL) model for the parallel configuration, exhibiting a higher value at T= 0 K as suggested by the dashed line in Figure 4D. The disparity between the in-plane and out-of-plane upper critical fields confirms the anisotropic nature of the superconductivity. Such anisotropy has been reported in various two-dimensional superconducting systems, including Nd1-xEuxNiO2, infinite-layer lanthanum nickelates, LaMnO3/KTaO3, and NbSe2 [36, 37, 17, 38].
Notably, the extrapolated in-plane upper critical field exceeds the Pauli paramagnetic limit, which is given by [39, 40]. Recent studies have demonstrated that strong spin-orbit coupling (SOC) in two-dimensional superconductors, can lead to out-of-plane spin locking and the formation of a uniaxial Ising spin texture[35, 41]. This spin-momentum locking suppresses Zeeman depairing, a key characteristic of Ising superconductivity, and can result in highly anisotropic superconducting behavior, with in-plane upper critical fields exceeding the Pauli limit. In this context, the enhanced in-plane upper critical field observed may be indicative of similar underlying physics.
Figure 5A presents the current-voltage (I-V) characteristic curves as a function of temperature for Eu/p-BBO sample (Sample 2). From the I-V measurements, the critical currents () are determined and plotted in Figure 5B. The decreases with increasing temperature and eventually reaches zero at the transition temperature, as expected for a superconductor. We observe that the slope of I-V characteristics transitions from a normal ohmic state () to a non-linear power law () at the superconducting transition temperature. In 2D superconductors, this transition can be understood in the framework of the Berezinskii–Kosterlitz–Thouless (BKT) transition, where the emergence of nonlinear I-V characteristics in the superconducting state results from the current-driven unbinding of vortex-antivortex pairs, which are generated by thermal fluctuations at finite temperatures [42, 43]. The I-V characteristics are also plotted on a logarithmic scale in Figure 5C, highlighting the non-linearity more distinctly. To further analyze this transition, we extracted the exponent from the I-V curves and plot as a function of temperature in Figure 5D. The BKT transition temperature () is estimated to be around 5.34 K by interpolating to , which closely matches obtained from resistance measurements. Moreover, the temperature-dependent resistance near TBKT is expected to follow the Halperin-Nelson (HN) formula: R = R0 exp(-bt-1/2 ), where R0 and b are material-specific parameters and t = T/TBKT - 1 [44]. Figure S3 presents the fit based on the HN formula, yielding TBKT = 5.31 K, which agrees well with the value obtained from the I-V analysis. This agreement with BKT theory provides strong evidence for the two-dimensional (2D) nature of superconductivity in the heterostructures.
Based on the structural transitions evidenced by the XRD and STEM measurements, the emergent superconductivity in the nominal Eu/p-BBO heterostructures arises from the formation of BaBi3 via the reduction of p-BBO. The structural results presented here rule out an electronic reconstruction at the interface between EuO and oxygen-deficient perovskite BaBiO3-x and/or electron doping of BaBiO3-x as the origin of superconductivity in nominal Eu/p-BBO heterostructures[20]. Studies of oxygen-deficient perovskite BaBiO3-x find an enhanced electronic band gap, suggesting that electron doping via oxygen vacancies may not stabilize superconductivity [45, 46]. We also consider the possibility of the reduction of p-BBO leading to the formation of an infinite-layered bismuthate structure analogous to the superconducting infinite-layer structures, which are also stabilized by the reduction of the perovskite phase[47, 48, 36, 49, 50]. First-principle calculations for an infinite-layer type BaBiO2 have recently been reported with a predicted Tc of 3.9–7.7 K under pressure and with spin-orbit coupling[51]. Here, superconductivity emerges on suppression of rotation of the oxygen sublattice at 10 GPa. However, the lattice parameters of the proposed structure are incompatible with our experimental results. It is, however, plausible that the layered infinite-layer phase forms as an intermediate phase and decomposes into BaO and BaBa3. For example, bulk stoichiometric NdNiO3 is known to be thermodynamically unstable and decomposes into Nd2O3, NiO, and Ni[52].
Conclusion
In conclusion, we observe quasi-two-dimensional superconductivity at 6 K in in situ reduced perovskite BaBiO3 films. The structural characterization of the superconducting films points to the decomposition of the reduced perovskite phase to form the superconducting BaBi3 phase. Although BaBi3 has been studied in bulk crystals due to the role of spin-orbit coupling in enhancing superconductivity, realizing superconductivity in films with reduced dimensionality allows investigating the effects of confinement and proximity effects with other bismuth-based layered topological quantum materials[53, 54].
II Materials and Methods
Sample growth
Thin films of perovskite-phase BaBiO3 were grown to thickness of 60-90 unit cells (uc) on 5 mm × 5 mm SrTiO3 (001) substrates (commercially available) using molecular beam epitaxy (MBE). Before thin film synthesis, the substrates were cleaned at 575oC in an activated radio frequency oxygen plasma at a chamber pressure of 6x10-6 Torr for 15 minutes. Bi, Ba, Eu and Al were evaporated from effusion cells, with their growth rates calibrated using a quartz crystal monitor. The p-BBO thin films were grown under the same conditions as the plasma cleaning process. After growth, the films were cooled in the oxygen plasma to 500∘C. Subsequently, the oxygen source was turned off, and 20-80 ML Eu metal were deposited in vacuum, followed by an additional 5-10 ML of Eu in an oxygen partial pressure of 2x10-9 Torr, as measured by a residual gas analyzer. Next, 10 ML of EuO was deposited in an oxygen partial pressure of 1x10-9 Torr. The films were then cooled to room temperature for the in situ deposition of Al, forming an AlOx capping layer. Finally, the samples were transferred for ex-situ X-ray diffraction and transport measurements.
Sample characterization
X-ray diffraction (XRD) and X-ray reflectivity (XRR) measurements were performed using a high-resolution X-ray diffractometer (Panalytical X’Pert Pro) equipped with a Ge (220) monochromator. The incident X-ray energy was fixed at the Cu-K wavelength of 8.04 keV. Temperature-dependent resistance and magnetotransport measurements were conducted using a Quantum Design Dynacool Physical Property Measurement System (PPMS) in the van der Pauw configuration. Al wires were ultrasonically bonded to the corners of the 5 mm × 5 mm samples to ensure reliable electrical contacts. The superconducting onset temperature () is defined as the temperature at which the resistance drops to 90% of the normal state resistance, while the superconducting zero temperature (Tc,0) is defined as the temperature at which the resistance reaches zero.
Cross-sectional STEM samples of the films were prepared by using a Ga+ ion beam in a Thermo Scientific Scios 2 DualBeam system. A probe aberration corrected Thermo Fisher Titan 80-300 kV microscope operating at 200 kV was used for STEM HAADF imaging and EDX measurements.
Acknowledgements.
D.P.K. and S. acknowledge support from the Duke Endowment Duke Science and Technology Fund. Transport measurements by J.T.M. and G.F. were supported by the National Science Foundation grant DMR-2327535. STEM imaging was performed in part at the Analytical Instrumentation Facility (AIF) at North Carolina State University, which is supported by the State of North Carolina and the National Science Foundation (award number ECCS-2025064). The AIF is a member of the North Carolina Research Triangle Nanotechnology Network (RTNN), a site in the National Nanotechnology Coordinated Infrastructure (NNCI).References
- [1] A Wl Sleight, JL Gillson, and PE Bierstedt. High-temperature superconductivity in the BaPb1-xBixO3 systems. Solid State Communications, 17(1):27–28, 1975.
- [2] LF Mattheiss, EM Gyorgy, and DW Johnson Jr. Superconductivity above 20 K in the Ba-K-Bi-O system. Physical Review B, 37(7):3745, 1988.
- [3] RJ Cava, Ba Batlogg, JJ Krajewski, Re Farrow, LW Rupp Jr, AE White, K Short, WF Peck, and T Kometani. Superconductivity near 30 k without copper: the Ba0.6K0.4BiO3 perovskite. Nature, 332(6167):814–816, 1988.
- [4] DG Hinks, B Dabrowski, JD Jorgensen, AW Mitchell, DR Richards, Shiyou Pei, and Donglu Shi. Synthesis, structure and superconductivity in the Ba1-x KxBiO3-y system. Nature, 333(6176):836–838, 1988.
- [5] Shiyou Pei, JD Jorgensen, B Dabrowski, DG Hinks, DR Richards, AW Mitchell, JM Newsam, SK Sinha, D Vaknin, and AJ Jacobson. Structural phase diagram of the Ba1-xKxBiO3 system. Physical Review B, 41(7):4126, 1990.
- [6] DE Cox and AW Sleight. Crystal structure of Ba2Bi3+Bi5+O6. Solid State Communications, 19(10):969–973, 1976.
- [7] Arthur W Sleight. Bismuthates: BaBiO3 and related superconducting phases. Physica C: Superconductivity and its Applications, 514:152–165, 2015.
- [8] Kateryna Foyevtsova, Arash Khazraie, Ilya Elfimov, and George A Sawatzky. Hybridization effects and bond disproportionation in the bismuth perovskites. Physical Review B, 91(12):121114, 2015.
- [9] DF Shao, X Luo, WJ Lu, L Hu, XD Zhu, WH Song, XB Zhu, and YP Sun. Spin-orbit coupling enhanced superconductivity in bi-rich compounds (A= Sr and Ba). Scientific Reports, 6(1):21484, 2016.
- [10] Lewis Powell, Wenjun Kuang, Gabriel Hawkins-Pottier, Rashid Jalil, John Birkbeck, Ziyi Jiang, Minsoo Kim, Yichao Zou, Sofiia Komrakova, Sarah Haigh, et al. Multiphase superconductivity in PdBi2. Nature Communications, 16(1):291, 2025.
- [11] BT Matthias and JK Hulm. A search for new superconducting compounds. Physical Review, 87(5):799, 1952.
- [12] Neel Haldolaarachchige, SK Kushwaha, Quinn Gibson, and RJ Cava. Superconducting properties of BaBi3. Superconductor Science and Technology, 27(10):105001, 2014.
- [13] Rajveer Jha, Marcos A Avila, and Raquel A Ribeiro. Hydrostatic pressure effect on the superconducting properties of BaBi3 and SrBi3 single crystals. Superconductor Science and Technology, 30(2):025015, 2016.
- [14] Yanan Wang, Tomoya Taguchi, Huan Li, Ai Suzuki, Yanting Zhang, Akari Miura, Mitsuki Ikeda, Hidenori Goto, Ritsuko Eguchi, Takafumi Miyazaki, et al. Superconducting properties of BaBi3 at ambient and high pressures. Physical Chemistry Chemical Physics, 23(40):23014–23023, 2021.
- [15] N Haberkorn, Raquel A Ribeiro, Li Xiang, SL Bud’ko, and Paul C Canfield. Superconducting properties and vortex pinning in intermetallic BaBi3 single crystals: a magnetization study. Physica C: Superconductivity and its Applications, 615:1354387, 2023.
- [16] Wen-Lin Wang, Yi-Min Zhang, Nan-Nan Luo, Jia-Qi Fan, Chong Liu, Zi-Yuan Dou, Lili Wang, Wei Li, Ke He, Can-Li Song, et al. Surface symmetry breaking and disorder effects on superconductivity in perovskite BaBi3 epitaxial films. Physical Review B, 98(6):064511, 2018.
- [17] Haoxiong Zhang, Awabaikeli Rousuli, Kenan Zhang, Laipeng Luo, Chenguang Guo, Xin Cong, Zuzhang Lin, Changhua Bao, Hongyun Zhang, Shengnan Xu, et al. Tailored ising superconductivity in intercalated bulk NbSe2. Nature Physics, 18(12):1425–1430, 2022.
- [18] Gang Li, Binghai Yan, Ronny Thomale, and Werner Hanke. Topological nature and the multiple dirac cones hidden in bismuth high-Tc superconductors. Scientific Reports, 5(1):10435, 2015.
- [19] Binghai Yan, Martin Jansen, and Claudia Felser. A large-energy-gap oxide topological insulator based on the superconductor BaBiO3. Nature Physics, 9(11):709–711, 2013.
- [20] Weiliang Qiao, Jiali Zhao, Yuxiao Chen, Shenggen Cao, Wenyu Xing, Ranran Cai, Liangliang Guo, Tian Qian, XC Xie, and Wei Han. Observation of quasi-two-dimensional superconductivity at the EuO-BaBiO3 interface. Physical Review B, 109(5):054509, 2024.
- [21] Wei Guo, Agham B Posadas, Sirong Lu, David J Smith, and Alexander A Demkov. Euo epitaxy by oxygen scavenging on SrTiO3 (001): Effect of SrTiO3 thickness and temperature. Journal of Applied Physics, 124(23):235301, 2018.
- [22] Thomas Mairoser, Julia A Mundy, Alexander Melville, Daniel Hodash, Paul Cueva, Rainer Held, Artur Glavic, Jürgen Schubert, David A Muller, Darrell G Schlom, et al. High-quality EuO thin films the easy way via topotactic transformation. Nature Communications, 6(1):7716, 2015.
- [23] Robert J Zollweg. X-ray lattice constant of barium oxide. Physical Review, 100(2):671, 1955.
- [24] BT Matthias, RM Bozorth, and JH Van Vleck. Ferromagnetic interaction in EuO. Physical Review Letters, 7(5):160, 1961.
- [25] Jie Yuan, Qihong Chen, Kun Jiang, Zhongpei Feng, Zefeng Lin, Heshan Yu, Ge He, Jinsong Zhang, Xingyu Jiang, Xu Zhang, Yujun Shi, Yanmin Zhang, Mingyang Qin, Zhi Gang Cheng, Nobumichi Tamura, Yi-feng Yang, Tao Xiang, Jiangping Hu, Ichiro Takeuchi, Kui Jin, and Zhongxian Zhao. Scaling of the strange-metal scattering in unconventional superconductors. Nature, 602:431-436, 2022.
- [26] Alekseĭ Alekseevich Abrikosov. Fundamentals of the Theory of Metals. NorthHolland, Amsterdam, 1988.
- [27] HM Jaeger, DB Haviland, BG Orr, and AM Goldman. Onset of superconductivity in ultrathin granular metal films. Physical Review B, 40(1):182, 1989.
- [28] Patrick Lömker and Martina Müller. Redox-controlled epitaxy and magnetism of oxide heterointerfaces: EuO/SrTiO3. Physical Review Materials, 3(6):061401, 2019.
- [29] Kristy J Kormondy, Lingyuan Gao, Xiang Li, Sirong Lu, Agham B Posadas, Shida Shen, Maxim Tsoi, Martha R McCartney, David J Smith, Jianshi Zhou, et al. Large positive linear magnetoresistance in the two-dimensional t2g electron gas at the EuO/SrTiO3 interface. Scientific Reports, 8(1):7721, 2018.
- [30] K Zou, Sohrab Ismail-Beigi, Kim Kisslinger, Xuan Shen, Dong Su, FJ Walker, and CH Ahn. LaTiO3/KTaO3 interfaces: A new two-dimensional electron gas system. APL Materials, 3(3), 2015.
- [31] Hongrui Zhang, Yu Yun, Xuejing Zhang, Hui Zhang, Yang Ma, Xi Yan, Fei Wang, Gang Li, Rui Li, Tahira Khan, et al. High-mobility spin-polarized two-dimensional electron gases at EuO/KTaO3 interfaces. Physical Review Letters, 121(11):116803, 2018.
- [32] Nand Kumar, Neha Wadehra, Ruchi Tomar, Shama, Sanjeev Kumar, Yogesh Singh, Sushanta Dattagupta, and Suvankar Chakraverty. Observation of shubnikov-de haas oscillations, planar hall effect, and anisotropic magnetoresistance at the conducting interface of EuO–KTaO3. Advanced Quantum Technologies, 4(1):2000081, 2021.
- [33] Changjiang Liu, Xi Yan, Dafei Jin, Yang Ma, Haw-Wen Hsiao, Yulin Lin, Terence M Bretz-Sullivan, Xianjing Zhou, John Pearson, Brandon Fisher, et al. Two-dimensional superconductivity and anisotropic transport at KTaO3 (111) interfaces. Science, 371(6530):716–721, 2021.
- [34] M Tinkham. Introduction to superconductivity, Mac-Graw-Hill Inc. New York, London, Tokyo, 1996.
- [35] JM Lu, O Zheliuk, Inge Leermakers, Noah FQ Yuan, Uli Zeitler, Kam Tuen Law, and JT Ye. Evidence for two-dimensional ising superconductivity in gated MoS2. Science, 350(6266):1353–1357, 2015.
- [36] Wenzheng Wei, Dung Vu, Zhan Zhang, Frederick J Walker, and Charles H Ahn. Superconducting Nd1-xEuxNiO2 thin films using in situ synthesis. Science Advances, 9(27):eadh3327, 2023.
- [37] Athby H Al-Tawhid, Samuel J Poage, Salva Salmani-Rezaie, Antonio Gonzalez, Shalinee Chikara, David A Muller, Divine P Kumah, Maria N Gastiasoro, José Lorenzana, and Kaveh Ahadi. Enhanced critical field of superconductivity at an oxide interface. Nano Letters, 23(15):6944–6950, 2023.
- [38] Wenjie Sun, Yueying Li, Ruxin Liu, Jiangfeng Yang, Jiayi Li, Wei Wei, Gangjian Jin, Shengjun Yan, Haoying Sun, Wei Guo, et al. Evidence for anisotropic superconductivity beyond pauli limit in infinite-layer lanthanum nickelates. Advanced Materials, 35(32):2303400, 2023.
- [39] BS Chandrasekhar. A note on the maximum critical field of high-field superconductors. Appl. Phys. Letters, 1:7–8, 1962.
- [40] Albert M Clogston. Upper limit for the critical field in hard superconductors. Physical Review Letters, 9(6):266, 1962.
- [41] Yu Saito, Yasuharu Nakamura, Mohammad Saeed Bahramy, Yoshimitsu Kohama, Jianting Ye, Yuichi Kasahara, Yuji Nakagawa, Masaru Onga, Masashi Tokunaga, Tsutomu Nojima, et al. Superconductivity protected by spin–valley locking in ion-gated MoS2. Nature Physics, 12(2):144–149, 2016.
- [42] MR Beasley, JE Mooij, and TP Orlando. Possibility of vortex-antivortex pair dissociation in two-dimensional superconductors. Physical Review Letters, 42(17):1165, 1979.
- [43] K Epstein, AM Goldman, and AM Kadin. Vortex-antivortex pair dissociation in two-dimensional superconductors. Physical Review Letters, 47(7):534, 1981.
- [44] B. I. Halperin and David R. Nelson. Resistive transition in superconducting films. J Low Temp Phys, 36:599-616, 1979.
- [45] Hui Cao, Hongli Guo, Yu-Cheng Shao, Qixin Liu, Xuefei Feng, Qinwen Lu, Zhongping Wang, Aidi Zhao, Atsushi Fujimori, Yi-De Chuang, et al. Realization of electron antidoping by modulating the breathing distortion in BaBiO3. Nano Letters, 21(9):3981–3988, 2021.
- [46] P Lightfoot, Joe A Hriljac, Shiyou Pei, Ying Zheng, AW Mitchell, DR Richards, B Dabrowski, JD Jorgensen, and DG Hinks. BaBiO2.5, a new bismuth oxide with a layered structure. Journal of Solid State Chemistry, 92(2):473–479, 1991.
- [47] Danfeng Li, Kyuho Lee, Bai Yang Wang, Motoki Osada, Samuel Crossley, Hye Ryoung Lee, Yi Cui, Yasuyuki Hikita, and Harold Y Hwang. Superconductivity in an infinite-layer nickelate. Nature, 572(7771):624–627, 2019.
- [48] Matthias Hepting, Danfeng Li, CJ Jia, Haiyu Lu, E Paris, Y Tseng, X Feng, M Osada, E Been, Y Hikita, et al. Electronic structure of the parent compound of superconducting infinite-layer nickelates. Nature Materials, 19(4):381–385, 2020.
- [49] Shengwei Zeng, Changjian Li, Lin Er Chow, Yu Cao, Zhaoting Zhang, Chi Sin Tang, Xinmao Yin, Zhi Shiuh Lim, Junxiong Hu, Ping Yang, et al. Superconductivity in infinite-layer nickelate La1-xCaxNiO2 thin films. Science Advances, 8(7):eabl9927, 2022.
- [50] Yusuke Nomura and Ryotaro Arita. Superconductivity in infinite-layer nickelates. Reports on Progress in Physics, 85(5):052501, 2022.
- [51] Yaning Cui, Heng Gao, Yongchang Li, Shaowen Xu, Haotian Wang, and Wei Ren. Spin-orbit coupling enhanced electron-phonon superconductivity in infinite-layer BaBiO2. AIP Advances, 13(12), 2023.
- [52] Oleksandr I Malyi, Julien Varignon, and Alex Zunger. Bulk ndnio2 is thermodynamically unstable with respect to decomposition while hydrogenation reduces the instability and transforms it from metal to insulator. Physical Review B, 105(1):014106, 2022.
- [53] Yuqi Xia, Dong Qian, David Hsieh, L Wray, Arijeet Pal, Hsin Lin, Arun Bansil, DHYS Grauer, Yew San Hor, Robert Joseph Cava, et al. Observation of a large-gap topological-insulator class with a single dirac cone on the surface. Nature Physics, 5(6):398–402, 2009.
- [54] Shuichi Murakami. Quantum spin hall effect and enhanced magnetic response by spin-orbit coupling. Physical Review Letters, 97(23):236805, 2006.